OpenStudy (anonymous):

you roll a 6 sided die with faces numbered 1 through 6, and toss a coin. what is the probability of rolling a 5 or getting tails?

OpenStudy (johnweldon1993):

Since the 2 are completely independent of each other... The probability of rolling a 5 is 1/6 right? And the probability of getting tails is 1/2 So the probability of both would be \[\large \frac{1}{6} \times \frac{1}{2} = ?\]

OpenStudy (anonymous):

so 2/12?

OpenStudy (johnweldon1993):

No, remember it is multiplication....1 times 1 is 1 :) \[\large \frac{1}{12}\]

OpenStudy (anonymous):

oh yeah! oops my bad haha. but thank you so much! you're a lifesaver.

OpenStudy (johnweldon1993):

lol no problem :)